Does anyone know what is the Hindus Formula for Quadratic Eqaution??
If yes please tell and state it and eplain?!

Grade:8

3 Answers

Ravi
9 years ago
It is the same quadratic formula that is often called by different names such as Sridharacharya’s formula, SDC formula or Hindu formula.
The roots of a quaratic eqn can be achieved a formula that accounts for real, negative as well as imaginary roots. It’s just that in the HIndu’s formula when it was derived in olden times, the concept of complex numbers was not available and hence, if the square of roots were negative, they were discarded. The roots for the quadratic eqn ax2+bx+c=0 is given by

x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}
Mohammed Anas
9 years ago
Thank You!
Prajwal Kavad
8 years ago
x=-b+OR- root over b^2-4ac/2a